Calculate the amount of heat needed to transform, under standard conditions, 1.0 mol of liquid methanol at 273 K to 1.0 mol of methanol vapor at 373 K. Assume that Cp,m for liquid methanol and for methanol vapor are independent of temperature. (Enthalpy of vaporization of methanol 35.3 KJ/mol-1, Cp,m of methanol 81.6 JK-1/mol)
Answer #1

q1 = no of moles*Cp*\DeltaT

    = 1*81.6*(373-273)

    = 8160joules   = 8.61KJ

q2 = no of moles*Enthalpy of vaporization of methanol

     = 1*35.3 KJ/mol    = 35.3KJ

q   = q1 +q2

    = 8.61 + 35.3

    = 43.91Kj >>>>> answer